OpenStudy (anonymous):

Find the area of the sector of the disk of radius r = 20 ft and with arc length s = 17 ft. Answer in units of ft^2

OpenStudy (anonymous):

Total area =pi* r^2 area of sector having arc length 17 ft and radius 20ft =(17/2 * pi*20)*pi * 20^2 =170 ft^2

OpenStudy (ybarrap):

Area = (s/(2*pi*r)* pi*r^2 = s*r/2 = 17*20/2 ft^2 = 17*10 ft^2 Ans. Area = 170 ft^2